# Power Optimization and Deep Learning for Channel Estimation of Active IRS-Aided IoT

### v1

@@ I-A Prior Works
Recently, some channel estimation (CE) schemes have been designed for IRS-assisted wireless networks [24, 25, 26]. In general, the CSI for direct channels (DCs) can be obtained through some traditional classical algorithms [27]. Some researchers have pointed out that the cascaded CE issue can be addressed through algorithms, such as least square (LS), or minimum mean square error (MMSE). It is worth noting that the construction of cascaded channels (CCs) involves the product of the number of base station (BS) antennas and the number of IRS elements. Specifically, compared to conventional large-scale multiple input multiple output (MIMO) CE schemes that do not include IRS, IRS assisted communication systems face the challenge of rapidly increasing pilot overhead in the cascaded CE process. To effectively mitigate the above mentioned pilot overhead, the authors of [28] designed a CE algorithm based on a three-phase pilot design, which was customized for IRS-assisted uplink multiuser communication systems. Furthermore, a low pilot overhead strategy for MIMO uplink transmission was explored in [29]. This study adopted a two-stage CE framework: in the first phase, an efficient pilot transmission scheme in time division duplex mode was used to focus on CE between BS and IRS. In the second phase, the BS-IRS channel information obtained in the first stage was utilized to further derive the channel status between IRS and IoT devices. This method not only improved the efficiency of CE, but also promoted the practical application of IRS assisted communication systems in resource limited scenarios.
@@ I-B Our Contributions
    Firstly, an active IRS-aided uplink IoT network is constructed. Unlike previous studies that often overlooked the estimation of DCs, the direct and cascaded channels are perfectly separated by adopting a clever pilot pattern. Subsequently, LS channel estimators for the DC and the CC are proposed, respectively. Correspondingly, the closed-form expressions of the MSE for both DC and CC are derived.
@@ I-C Organization and Notation
The remainder of this paper is organized as follows. In Section II, the system model of an active IRS-aided uplink IoT network is constructed. In Section III, we first propose the LS channel estimator and MMSE estimator, then, the optimal PA strategy is investigated to minimize Sum-MSE. In Section IV, the CDCE algorithm and the CCCE algorithm are proposed, and the computational complexity of the two proposed networks is also given. Simulation results and conclusions are provided in Section V and Section VI, respectively.
Notations: Throughout the paper, vectors and matrices are denoted by boldface lower case and upper case letters, respectively. Lower case letters are adopted to represent scalars, while $(\cdot)^{H}$ stands for conjugate and transpose operation. The sign $(\cdot)^{T}$ denotes the transpose operation. The signs $\|\cdot\|_{2}$ and $\|\cdot\|_{F}$ stand for the 2-norm and $F$-norm, respectively. The notation $\mathbb{E}\{\cdot\}$ denotes the expectation operation. $\hat{[\;]}$ represents the estimation operation.
@@ II System model
The active IRS-assisted uplink IoT network is shown in Fig. 1, where the IoT device is configured with single antenna, the BS is equipped with $K$ antennas, and IRS employs $N$ active reflecting elements. It is worth noting that compared to passive IRS that can only passively reflect signals, active IRS is unique due to the integration of reflection-type amplifiers inside its elements. In other words, active IRS not only actively reflects signals, but also has the ability to amplify signals. Without loss of generality, it is assumed that all channels are Rayleigh fading channels. Due to the property of channel reciprocity, the CSI of the downlink can be obtained by estimating the uplink channel.
@@
where $x$ is the transmit symbol with unit energy, $P_{t}$ represents the sum of the transmit power at IoT device and the power reflected by active IRS, and $\beta$ denotes the PAF.
@@ III-A Proposed LS Channel Estimator
where the number of columns in matrix $\mathbf{V}$ is chosen to be greater than or equal to the number of its rows to guarantee the existence of the right pseudo-inverse of $\mathbf{V}$, i.e., $Q\geq N$. To minimize overheads, it is assumed that $Q=N$. Subsequently, by right multiplying $\mathbf{x}^{*}$ on both sides of (12) at the same time, we can yield
@@
Similarly, the MMSE estimator of the CC can be represented as follows
@@ IV-A Proposed CDCE algorithm
where $T$ represents the size of training dataset. Moreover, the conventional LS CE scheme can be employed to acquire the label $\mathbf{h}_{\text{LS}}$. The objective of training the CDCE network is to minimize $\mathcal{L}(\boldsymbol{\theta})$ by optimizing $\boldsymbol{\theta}$, which is
@@
where $\mathbf{g}(\boldsymbol{\theta}_{i})$ is the gradient vector for $\boldsymbol{\theta}_{i}$, and $\alpha_{i}$ is the learning rate. After the proposed CDCE network is trained, the BS can directly estimate the DC based on the well-trained CDCE network.
As displayed in Fig. 3, the proposed CDCE network consists of three convolution layers, three max pooling layers, and one linear layer. The spatial features of the channel matrix are exploited by combining the convolution (Conv) operation and the rectified linear unit (ReLU). To enhance the stability of the network and expedite the training speed of the proposed CDCE network, a batch normalization (BN) is introduced between Conv and ReLU. The first, third, and fifth layers are both the “Conv + BN + ReLU” operations. As $\tilde{\mathbf{Y}}_{1}$ is a complex-valued, it is initially crucial to extract the real part and the imaginary part of $\tilde{\mathbf{Y}}_{1}$ separately as the two input feature maps of the first convolution layer. This operation successfully converts the input of the proposed CDCE network into the real-valued. The second, fourth, and sixth layers are the max pooling layer, where the kernel size and the stride are set to 2. By deploying pooling layers to reduce the dimensionality of features extracted from previous layers, the parameter and computational complexity of the next layer are reduced, significantly reducing the consumption of computing resources while ensuring model performance. Additionally, pooling layers prevent overfitting and render the features more robust against noise, enhancing the overall stability and generalization capabilities of the model. The seventh layer is a linear layer, which maps the output of the previous layer into a vector with $K\times 1\times 2$ elements, where the first $K$ elements denote the real part of $\hat{\mathbf{h}}_{\text{CDCE}}$, and the last $K$ elements correspond to its imaginary part. The hyperparameters of the proposed CDCE network for DC are depicted in Table I.
@@ 7 Linear $(\frac{K}{8}\times\frac{N}{8}\times 128)\times(K\times 1\times 2)$
The proposed CDCE algorithm consists of two parts: training and testing, as shown in Algorithm 1, where $i$ and $I$ denote the iteration index and the maximum iteration number, respectively.
@@ IV-B Proposed CCCE algorithm
where $\beta_{i}$ is the learning rate.
@@ IV-C Computational Complexity Analysis
To further analyze the computational complexity of the proposed CDCE / CCCE network, it is first necessary to know the computational complexity of the convolutional layer, linear layer, and max pooling layer.
For a convolution layer, its computational complexity [40] can be expressed by
@@ IV-C1 For the proposed CDCE network
The computational complexity of the first convolution layer is
@@ IV-C2 For the proposed CCCE network
Based on Fig. 4 and Table II, the computational complexity of the first, third, and fifth convolutional layers is
@@
In addition, the computational complexity of the 7th, 8th, and 9th convolution layer is the same, which is
@@ V Simulation Results
In this section, firstly, the simulation results of the optimal PA strategy that can minimize Sum-MSE are provided. Subsequently, we present the simulation results to validate the performance and effectiveness of the proposed CDCE / CCCE network.
The distances from IoT devices to BS, from IoT devices to IRS, and from IRS to BS are 100 m, 70 m, and 60 m, respectively. The path loss at distance d is modeled as $\text{PL}(d)=\text{PL}_{0}-10\alpha\log_{10}(\frac{d}{d_{0}})$, where $\text{PL}_{0}$ = -30 dB stands for the path loss at a reference distance of $d_{0}$ = 1m, and $\alpha$ represents the path loss exponent. Specifically, the path loss exponents of the IoT devices $\to$ BS, IoT devices $\to$ IRS and IRS $\to$ BS are set as 3.5, 2.4, and 2.2, respectively.
Assuming 10000 samples are collected at BS, we proposed that 90% be allocated to form the training dataset, while the remaining 10% be designated for the validation dataset. In addition, we regenerate the 5000 samples used to appraise the performance of the proposed CDCE / CCCE network. Moreover, the AdamW optimization algorithm is employed to update the model’s weight parameters. The training process encompasses a total of 200 epochs. After 200 epochs, we will use the trained network to evaluate its performance. The learning rates of the proposed CDCE and CCCE networks are initialized to 1e-3 and 2.5e-4, respectively, and then reduced to the half of the original level every 35 epochs.
@@
Fig. 6 depicts the Sum-MSE versus the PAF $\beta$ with different number of BS antennas $K$. It is evident from the figure that as $K$ increases, the optimal PAF $\beta^{\text{opt}}$ decreases. This indicates that the power allocated to the BS decreases, while the power allocated to the active IRS increases. In addition, as $K$ increases, the value of Sum-MSE also exhibits an upward trend.
@@
Fig. 7 illustrates the Sum-MSE versus the PAF $\beta$ with different number of active IRS elements $N$. It is clear that as $N$ increases, the value of Sum-MSE decreases, which is consistent with the derived in (24). Furthermore, $\beta$ decreases as $N$ increases. This indicates that a larger proportion of power needs to be allocated to the active IRS.
@@
Fig. 9 shows the derived closed-form expression of the optimal PAF in (III-B1) versus the number of active IRS elements $N$ with different number of BS antennas $K$. For a given $N$, the optimal PAF $\beta^{\text{opt}}$ becomes smaller as $K$ increases. This trend aligns with the observations made in Fig. 6. Moreover, when $K$ is held constant, the optimal PAF $\beta^{\text{opt}}$ also exhibits a decreasing trend with an increase in $N$. This particular trend is consistent with the findings presented in Fig. 7.
@@
The training and validation performance of the proposed CDCE network and CCCE network are depicted in Fig. 10. We initially adopted a high learning rate to expedite the model training process and accelerate parameter updates. Subsequently, the learning rate was decayed by half every 35 epochs to facilitate better convergence of the model. As shown in Fig. 10, as the number of epochs progresses in the training of our proposed model, the corresponding training and validation losses undergo a rapid decline in the initial phase and gradually stabilize in the later stages. Concurrently, our strategy of learning rate decay effectively contributes to the convergence of the model.
@@
Fig. 12 displays the MSE performance versus different number of BS antennas $K$ when the number of active IRS elements $N=128$. As $K$ increases, it is apparent that both the MSE of the DC and the MSE of the CC also increase. This phenomenon is mainly because the size of the estimated channel increases with $K$. Furthermore, the trend observed in Fig. 12 aligns with the conclusions drawn from Fig. 6.

### v2

@@ I-A Prior Works
Recently, some channel estimation (CE) schemes have been designed for IRS-assisted wireless networks [24, 25, 26]. In general, the CSI for direct channels (DCs) can be obtained through some traditional classical algorithms [27]. Some researchers have pointed out that the cascaded CE issue can be addressed through algorithms, such as least square (LS), or minimum mean square error (MMSE). It is worth noting that the construction of cascaded channels (CCs) involves the product of the number of base station (BS) antennas and the number of IRS elements. Specifically, compared to conventional large-scale MIMO CE schemes that do not include IRS, IRS assisted communication systems face the challenge of rapidly increasing pilot overhead in the cascaded CE process. To effectively mitigate the above mentioned pilot overhead, the authors of [28] designed a CE algorithm based on a three-phase pilot design, which was customized for IRS-assisted uplink multiuser communication systems. Furthermore, a low pilot overhead strategy for MIMO uplink transmission was explored in [29]. This study adopted a two-stage CE framework: in the first phase, an efficient pilot transmission scheme in time division duplex mode was used to focus on CE between BS and IRS. In phase II, the BS-IRS channel information obtained in the first stage was utilized to further derive the channel status between IRS and IoT devices. This method not only improved the efficiency of CE, but also promoted the practical application of IRS assisted communication systems in resource limited scenarios.
@@ I-B Our Contributions
    Firstly, an active IRS-aided uplink IoT network is constructed. Unlike previous studies that often overlooked the estimation of DCs, the DCs and CCs are perfectly separated by adopting a clever pilot pattern. Subsequently, LS channel estimators for the DC and the CC are proposed, respectively. Correspondingly, the closed-form expressions of the MSE for both DC and CC are derived.
@@ I-C Organization and Notation
The remainder of this paper is structured as follows. In Section II, the system model of an active IRS-aided uplink IoT network is established. In Section III, we first propose the LS channel estimator and MMSE estimator, then, the optimal PA strategy is investigated to minimize Sum-MSE. In Section IV, the CDCE algorithm and the CCCE algorithm are proposed, and the computational complexity of the two proposed networks is also given. Simulation results and conclusions are provided in Section V and Section VI, respectively.
Notations: Throughout the paper, vectors and matrices are denoted by boldface lower case and upper case letters, respectively. Lower case letters are employed to represent scalars, while $(\cdot)^{H}$ stands for conjugate and transpose operation. The sign $(\cdot)^{T}$ denotes the transpose operation. The signs $\|\cdot\|_{2}$ and $\|\cdot\|_{F}$ stand for the 2-norm and $F$-norm, respectively. The notation $\mathbb{E}\{\cdot\}$ denotes the expectation operation. $\hat{[\;]}$ represents the estimation operation.
@@ II System model
The active IRS-assisted uplink IoT network is shown in Fig. 1, where the IoT device is configured with single antenna, the BS is equipped with $K$ antennas, and IRS employs $N$ active reflecting elements. It is worth noting that compared to passive IRS that can only passively reflect signals, active IRS is unique due to the integration of reflection-type amplifiers inside its elements. In other words, active IRS not only actively reflects signals, but also has the ability to amplify signals. Assuming that IoT devices to active IRS, IoT devices to BS, and active IRS to BS are all Rayleigh fading channels. It is worth noting that BS can estimate the channel characteristics between a large number of antennas through a small number of pilot signals in the uplink. Based on channel reciprocity, it can effectively predict or reconstruct the CSI of the downlink.
@@
where $x$ is the transmit symbol with unit energy, $P_{t}$ represents the power sum of the IoT devices and active IRS, and $\beta$ denotes the PAF.
@@ III-A Proposed LS Channel Estimator
where $Q\geq N$ is used to ensure the existence of the right pseudo-inverse of $\mathbf{V}$. To minimize overheads, it is assumed that $Q=N$. Subsequently, by right multiplying $\mathbf{x}^{*}$ on both sides of (12) at the same time, we can yield
@@
Similarly, the MMSE estimator of the CC is
@@ IV-A Proposed CDCE algorithm
where $T$ represents the size of training dataset. Moreover, the conventional LS CE scheme can be employed to acquire the label $\mathbf{h}_{\text{LS}}$. The goal of training the CDCE network is to minimize $\mathcal{L}(\boldsymbol{\theta})$ by optimizing $\boldsymbol{\theta}$, which is
@@
where $\mathbf{g}(\boldsymbol{\theta}_{i})$ is the gradient vector (GV) for $\boldsymbol{\theta}_{i}$, and $\alpha_{i}$ is the learning rate (LR). After the proposed CDCE network is trained, the BS can directly estimate the DC based on the well-trained CDCE network.
As displayed in Fig. 3, the proposed CDCE network consists of three convolution layers, three max pooling layers, and one linear layer. The spatial features of the channel matrix are exploited by combining the convolution (Conv) operation and the rectified linear unit (ReLU). To enhance the stability of the network and expedite the training speed of the proposed CDCE network, a batch normalization (BN) is introduced between Conv and ReLU. The 1st, 3rd, and 5th layers are both the “Conv + BN + ReLU” operations. As $\tilde{\mathbf{Y}}_{1}$ is a complex-valued, it is initially crucial to extract the real part and the imaginary part of $\tilde{\mathbf{Y}}_{1}$ separately as the two input feature maps of the first convolution layer. This operation successfully converts the input of the proposed CDCE network into the real-valued. The second, fourth, and sixth layers are the max pooling layer, where the kernel size and the stride are set to 2. By deploying pooling layers to reduce the dimensionality of features extracted from previous layers, the parameter and computational complexity of the next layer are reduced, significantly reducing the consumption of computing resources while ensuring model performance. Additionally, pooling layers prevent overfitting and render the features more robust against noise, enhancing the overall stability and generalization capabilities of the model. The seventh layer is a linear layer, which maps the output of the previous layer into a vector with $K\times 1\times 2$ elements, where the first $K$ elements denote the real part of $\hat{\mathbf{h}}_{\text{CDCE}}$, and the last $K$ elements correspond to its imaginary part. The hyperparameters of the proposed CDCE network for DC are depicted in Table I.
@@ 7 Linear $(\frac{K}{8}\times\frac{N}{8}\times 128)\times(K\times 1\times 2)$
As presented in Algorithm 1, the proposed CDCE algorithm contains a training part and a testing part, where $i$ is the iteration index and $I$ denotes the maximum iteration number.
@@ IV-B Proposed CCCE algorithm
where $\mathbf{g}(\boldsymbol{\delta}_{i})$ is the GV for $\boldsymbol{\delta}_{i}$, and $\beta_{i}$ is the LR.
@@ IV-C Computational Complexity Analysis
For a Conv layer, its computational complexity [40] can be expressed by
@@ IV-C1 For the proposed CDCE network
The computational complexity of the 1st Conv layer is
@@ IV-C2 For the proposed CCCE network
Based on Fig. 4 and Table II, the computational complexity of the 1st, 3rd, and 5th Conv layers is
@@
In addition, the computational complexity of the 7th, 8th, and 9th Conv layer is the same, which is
@@ V Simulation Results
In this section, first and foremost, the simulation results of the optimal PA strategy capable of minimizing Sum-MSE are presented. Subsequently, the simulation results are provided to validate the performance and effectiveness of the proposed CDCE / CCCE network.
The distances from IoT devices to BS, from IoT devices to active IRS, and from active IRS to BS are 100 m, 70 m, and 60 m, respectively. The path loss at distance $d$ is modeled as follows
$\text{PL}(d)=\text{PL}_{0}-10\alpha\log_{10}(\frac{d}{d_{0}}),$ (65)
where $\text{PL}_{0}$ = -30 dB stands for the path loss at a reference distance of $d_{0}$ = 1m, and $\alpha$ indicates the path loss exponent. Specifically, the path loss exponents of the IoT devices $\to$ BS, IoT devices $\to$ active IRS and active IRS $\to$ BS are set as 3.5, 2.4, and 2.2, respectively.
Assuming 10000 samples are collected at BS, we proposed that 90% be allocated to form the training dataset, while the remaining 10% be designated for the validation dataset. In addition, we regenerate the 5000 samples used to appraise the performance of the proposed CDCE / CCCE network. Moreover, the AdamW optimization algorithm is employed to update the model’s weight parameters. The training process encompasses a total of 200 epochs. After 200 epochs, we will use the trained network to evaluate its performance. The LRs of the proposed CDCE and CCCE networks are initialized to 1e-3 and 2.5e-4, respectively, and then reduced to the half of the original level every 35 epochs.
@@
Fig. 6 depicts the Sum-MSE versus the PAF $\beta$ with different $K$. It is evident from the figure that as $K$ increases, the optimal PAF $\beta^{\text{opt}}$ decreases. This indicates that the power allocated to the BS decreases, while the power allocated to the active IRS increases. In addition, as $K$ increases, the value of Sum-MSE also exhibits an upward trend.
@@
Fig. 7 illustrates the Sum-MSE versus the PAF $\beta$ with different $N$. It is clear that as $N$ increases, the Sum-MSE decreases, which is consistent with the derived in (24). Furthermore, $\beta$ decreases as $N$ increases. This indicates that a larger proportion of power needs to be allocated to the active IRS.
@@
Fig. 9 shows the derived closed-form expression of the optimal PAF in (III-B1) versus $N$ with different $K$. For a given $N$, the optimal PAF $\beta^{\text{opt}}$ becomes smaller as $K$ increases. This trend aligns with the observations made in Fig. 6. Moreover, when $K$ is held constant, the optimal PAF $\beta^{\text{opt}}$ also exhibits a decreasing trend with an increase in $N$. This particular trend is consistent with the findings presented in Fig. 7.
@@
The training and validation performance of the proposed CDCE network and CCCE network are depicted in Fig. 10. We initially adopted a high LR to expedite the model training process and accelerate parameter updates. Subsequently, the LR was decayed by half every 35 epochs to facilitate better convergence of the model. As shown in Fig. 10, as the number of epochs progresses in the training of our proposed model, the corresponding training and validation losses undergo a rapid decline in the initial phase and gradually stabilize in the later stages. Concurrently, our strategy of LR decay effectively contributes to the convergence of the model.
@@
Fig. 12 displays the MSE performance versus different $K$ when $N=128$. As $K$ increases, it is apparent that both the MSE of the DC and the MSE of the CC also increase. This phenomenon is mainly because the size of the estimated channel increases with $K$. Furthermore, the trend observed in Fig. 12 aligns with the conclusions drawn from Fig. 6.